# RIVCoin: an alternative, integrated, CeFi/DeFi-Vaulted Cryptocurrency

### v1

@@ Abstract
This whitepaper introduces RIVCoin, a cryptocurrency built on Cøsmos, fully stabilized by a diversified portfolio of both CeFi and DeFi assets, available in a digital, non-custodial wallet called RIV Wallet, that aims to provide Users an easy way to access the cryptocurrency markets, compliant to the strictest AML laws and regulations up to date. The token is a cryptocurrency at any time stabilized by a basket of assets: reserves are invested in a portfolio composed long term by 50% of CeFi assets, comprised of Fixed Income, Equity, Mutual and Hedge Funds and 50% of diversified strategies focused on digital assets, mainly staking and LP farming on the major, battle tested DeFi protocols. The cryptocurrency, as well as the dollar before Bretton Woods, is always fully stabilized by vaulted proof of assets: it is born and managed as a decentralized token, minted by a Decentralized Autonomous Organization, and entirely stabilized by assets evaluated by professional independent third parties. Users will trade, pool, and exchange the token without any intermediary, being able to merge them into a Liquidity Pool whose rewards will be composed by both the trading fees and the liquidity rewards derived from the reserve’s seigniorage, that should affect the token’s price movement. In the long run, RIVCoin holders will also have access to an ecosystem of added-value services that will further increase the token’s value. Our cryptocurrency is built to be Proof of Stake (PoS - energy saver), Proof of Asset (PoA - stabilized) and Proof of Liquidity (PoL – market provided).
RIVCoin allows the User to enter the cryptocurrency market easily, without experiencing unjustified huge price depreciations, being the reserves pledged in last resort in favour of the users. Moreover, using RIV Wallet will allow the User to perform KYC/AML procedures that comply with the latest international regulatory FATF-GAFI VASP framework. The Liquidity Pool fair incentive mechanism is executed such that it will force a de facto democratic redistribution of wealth: Users who wish and decide to pool RIVCoin in the Liquidity Pool will receive additional RIVCoin for themselves, and new RIVCoin are minted when the reserves increase in value or in case of purchase of new RIVCoin. All wealthier Users are then accepting a redistribution of income, to the benefit of those who have purchased less tokens. In (Cooperative) Game Theory, maximization of the economic benefit of the ecosystem is achieved when players’ incentives are perfectly aligned. The proposed model allows for alignment of incentives: decreasing the risk exposure by wealthier Users, but implicitly increasing that of smaller ones to a level perceived by them as still sustainable and never creating ultra-speculative positions (according to H.P. Minsky definition, “when the incoming flows are not sufficient even to pay interest, so that it is necessary to apply for new loans both to repay the principal portion of the initial loan but also to honor the payment of the related interest”). In other words, wealthier Users stabilize the risk associated with the market portfolios of the reserves invested in Centralized and Decentralized Finance, without falling into the “bet scheme”. Users indirectly benefit from the access to the rewards of sophisticated cryptocurrency portfolios hitherto precluded to them, as well as having access to a real redistribution of wealth, without this turning into a disadvantage for the wealthy User, who benefits from the greater stability created by the huge influx of smaller Users. Therefore, the progressive growth becomes additional value that tends to stabilize over time, optimizing RIVCoin on the systemic risk level.
@@ 1 Introduction
The access of end Users to the cryptocurrency market is one of the most crucial points to solve before cryptocurrencies can reach mainstream adoption. Although cryptocurrency has seen a spike in interest in 2021, bolstered by the last Bitcoin halving occurred in May 2020, still mainstream adoption is far from happening.
Currently, Users access the markets by the means of centralized entities (usually via Centralized Exchanges), while only the more tech-savvy Users can quickly interact with Staking Pools, Liquidity Pools and all the various DeFi protocols. While using Centralized Exchanges is usually straightforward, still the User needs to trust a centralized entity, that often does not act in the best interest of its Users.
The aim of this whitepaper is to describe the first CeFi/DeFi-stabilized cryptocurrency built on Cøsmos and whose cash reserves are invested in a diversified combination of CeFi and DeFi assets with high degree of liquidity. The token, named RIVCoin, is completely guaranteed by its reserves, and has as its primary objective to allow everyone to easily access the cryptocurrency market via carefully enforced, self-imposed boundaries.
In the system we are designing, the User can only operate after proper KYC/AML procedures have been carried out: as a result, there is total transparency on who the Users are, drastically reducing the possibility of frauds, criminal or improper use of the token.
The problem of how retail Users access the cryptocurrency market is not to be overlooked. Trust in cryptocurrency slows at its minimum when scams happens. The recent cases of CEX bankruptcies have shed a bad light on crypto assets: while these have been isolated cases of mismanagement of the User’s funds, still the whole cryptocurrency market suffered from these crashes.
Crucial for every cryptocurrency is now to be, far beyond any other doubt, Proof of Stake (PoS - energy saver), Proof of Asset (PoA - stabilized) and Proof of Liquidity (PoL – market provided). These three core features contribute to the final tangible concept of cryptocurrency: its circulation as a medium of exchange adds value and does not represent a danger (i.e., energy-wise), it is accepted given its stabilized intrinsic value (constantly officially evaluated), it can be always exchanged either through the Liquidity Pools or through the Vault which will act always as the Seller of Last Resort (SoLR).
The is no direct linking between the RIVCoin value and the value of the underlying reserves. Our purpose is to create a cryptocurrency whose value should gradually increase towards a reference value, with possible fluctuations around such value. The reference value will change according to market conditions, while the primary minting price could temporarily be above or below such price, to stabilize the value of the currency.
In the system we are going to build, the Vault guarantees the total number of coins minted and, as a last resort, it can always be sold to reimburse part/all the coins in circulation. The Vault Reserves are invested in a portfolio of productive assets. They fluctuate in its value, but they cannot be used for anything else except for guaranteeing the reference value of all the coins minted.
The potential “bad inflation” is a-priori defined by the total number of coins mintable. The seignorage will be used for a democratic redistribution of the wealth generated by the reserves of the Vault, subject to explicit actions to be implemented by the Users. Like Central Banks redistribute the incomes to the national banks pro-rata their participation, the same will happen with our proposed model, which will redistribute the seigniorage to all the holders that actively operate as Liquidity Providers, as an extra incentive to the normal trading fees that every Liquidity Pool redistributes. Differently by what happens in the real economy, where the seigniorage increases with the monetization of public spending not covered by tax revenues, thus creating “bad inflation”, here the seigniorage is constantly redistributed and cannot be affected by new print of money. To some extent, it generates “good inflation”.
1.1 Towards a Decentralized Store of Value
Fiat-pegged stablecoins have emerged throughout the years as a solution to the high volatility of the crypto space. The need of having a numeraire which can hold value, being an effective mean of payment and unit of account, is of the utmost importance for cryptocurrencies to be considered actual currencies.
According to the Bank of International Settlements [2], stablecoins “evolved in order to address the failure of Bitcoin and other cryptocurrencies to provide an effective monetary and payment instrument. This reflected the preference of main market participants to base transactions and payments on sovereign fiat currencies, in particular the US dollar”.
It is evident that the need of a stable unit of account is essential for crypto to reach mainstream adoption. Still, the relatively small market capitalization, the macro-economic uncertainties and the low appetite for risk that nowadays characterize markets hinder this possibility from the start. The need for regulation is evident, now: the more the market capitalization of cryptocurrencies and digital assets grows, the more regulators will try to capture this unregulated market into a normative frame.
Central Bank Digital Currencies (CBDCs) have emerged as a possible stable unit of account in the cryptocurrency markets. Their potential to allow faster settlement of transactions and to reduce settlement risk is an interesting innovation in Banking and Financial Markets.
Many different regulators have taken initiatives towards the development of CBDCs [9]. Among the most important projects, the National Bank of Cambodia has launched in 2020 the first CBDC in the Asia region. Other important projects have been carried out by China and Singapore. Still, Central Bank Digital Currencies suffer from centralization: it is the Central Bank that issues the currency and guarantees for the value of the currency.
There have been many classifications of stablecoins in literature. The one of G. Liao and J.Caramichael [15] classifies stablecoins as follows:
@@ 1.1 Towards a Decentralized Store of Value
  • Public Reserve Backed Stablecoins, such as Tether, USDC and Paxos Dollar;

### v2

# RIVCoin_an_alternative_integrated_CeFi_DeFi_vaulted_cryptocurrency.bib

@@ Abstract
This whitepaper introduces RIV Coin, a cryptocurrency that is fully stabilized by a diversified portfolio of invested reserves that are a) evaluated by professional independent third parties, and b) auditable and provable by the protocol. It is born and managed as a decentralized token, minted by a Decentralized Autonomous Organization (DAO). All wealthier Users are then accepting a redistribution of income, to the benefit of those who have purchased less tokens. In cooperative Game Theory, maximization of the economic benefit of the ecosystem is achieved when players’ incentives are perfectly aligned. The proposed model allows for alignment of incentives: decreasing the risk exposure by wealthier Users, but implicitly increasing that of smaller ones to a level perceived by them as still sustainable and never creating ultra-speculative positions. In other words, wealthier Users stabilize the risk associated with the market value of portfolios in which the reserves are invested in Centralized and Decentralized Finance, without falling into the “bet scheme”. Users indirectly benefit from the access to the rewards of sophisticated cryptocurrency portfolios hitherto precluded to them, as well as having access to a real redistribution of wealth, without this turning into a disadvantage for the wealthy User, who benefits from the greater stability created by the huge influx of smaller Users. Therefore, the progressive growth becomes additional value that tends to stabilize over time, optimizing RIV Coin on the systemic risk level.
@@ 1 Introduction
[…]
The broader decentralized finance (DeFi) ecosystem has since struggled with this core challenge: reconciling decentralization with economic stability. While cryptocurrencies have innovated in removing custodial trust, they have not yet realized a system capable of adapting dynamically to economic realities without reintroducing centralized points of failure.
Meanwhile, traditional fiat currencies achieve relative stability through active monetary interventions–central banks managing interest rates and liquidity to stabilize economies. However, such systems are fundamentally dependent on perpetual debt expansion, introducing long-term risks of monetary devaluation and systemic fragility.
Despite their structural differences, both fiat and decentralized currencies ultimately derive value from trust: one rooted in institutions, the other in algorithms. Yet neither has fully solved the problem of creating a truly stable, decentralized medium of exchange anchored in real, productive value.
In the following sections, we explore the emergence of stablecoins as a bridge solution to these challenges, and examine why a more fundamental shift is needed.
1.1 Background
In response to the volatility of early cryptocurrencies, fiat-pegged stablecoins emerged as an attempt to introduce price stability into the digital asset ecosystem. By anchoring digital tokens to sovereign currencies like the US Dollar, stablecoins such as Tether and USDC sought to create a reliable medium of exchange within crypto markets. According to the Bank for International Settlements (BIS) in October 2019 in its “Investigating the impact of global stablecoins” [bis]
“The first wave of cryptoassets, of which Bitcoin is the best known, have so far failed to provide a reliable and attractive means of payment or store of value. This reflected the preference of main market participants to base transactions and payments on sovereign fiat currencies, particularly the US Dollar.”
Various taxonomies of stablecoins have been proposed in the literature. For instance, G. Liao and J. Carmichael [liao] classify stablecoins into:
@@ 1.1 Background
  • Public Reserve-Backed Stablecoins, such as Tether, USDC, and Paxos Dollar;